OpenStudy (anonymous):

Find the area of the indicated region. We suggest you graph the curves to check whether one is above the other or whether they cross, and that you use technology to check your answer. HINT [See Example 1.] Between y = x2 and y = −1 for x in [−1, 1]

OpenStudy (anonymous):

the area of the indicated region would be below the y=x^2 curve and above the y=-1 the area between the y=-1 line to the x axis is 1*2 = 2 units^2 the area below the y=x^2 to the x-axis is the ingeral of x^2 from -1 to 1. so, can you integrate the eqauation x^2 from -1 to 1? 

OpenStudy (anonymous):